OpenStudy (anonymous):

6+4x-4=8(2+x)-2(x-8) Show the steps

OpenStudy (saifoo.khan):

4x + 2 = 16 + 8x - 2x + 16 4x + 2 = 6x + 32

OpenStudy (saifoo.khan):

can you solve now?

OpenStudy (anonymous):

Yes thankyou. I have another 3k+3(6k-7)=8k+31